OpenStudy (anonymous):

There are 20 players on the school's water polo team. 8 of them are freshmen. The coach needs 8 starters. Determine the probability that the 8 freshmen will be randomly selected to be starters. A. 7/5,168 B. 1/125,970 C. 7/41,990 D. 1/5,168

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

wouldnt my answer be C?

OpenStudy (anonymous):

you know how to do this... give it a shot and I'll check your answer. oops, just a sec...

OpenStudy (anonymous):

not C

OpenStudy (anonymous):

it looks like A

OpenStudy (anonymous):

nope, looks like you're guessing...

OpenStudy (anonymous):

so think of it, at the beginning what is the probability to choose a freshman ?

OpenStudy (anonymous):

1 right?

OpenStudy (anonymous):

1/5,168

OpenStudy (anonymous):

look, you have a team of 20 players and 8 of them are freshman so the probability that the first one that you choose is a freshman is 8/20

OpenStudy (anonymous):

understand so far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, when we took one, we left with a team of 19 where 7 of them are fresh .. so what is the probability for the second one to be fresh ?

OpenStudy (anonymous):

12 right?

OpenStudy (anonymous):

why 12 ? (the probability cant be larger than 1 !!)

OpenStudy (anonymous):

okay im very confused atm, lol then obiously it must be 1 lol.

OpenStudy (anonymous):

look , when we first had 20 people and 8 of them were fresh the probability to choose fresh was 8/20

OpenStudy (anonymous):

now, we have 19 people and 7 of them are fresh so what is the probability to choose a fresh now ?

OpenStudy (anonymous):

it can also be thought of using combinations... there are 8 starters and 20 players. since you are choosing without replacement and order doesn't matter you can you combinations. the total number of different starting lineups are \[\left(\begin{matrix}20 \\ 8\end{matrix}\right)\] this is the denominator. Since there are 8 freshman and 12 non-freshmen and you need 8 of the freshmen and none of the non-freshmen, there is only 1 way to get this outcome: by choosing all of the freshmen. so the numerator would look like this: \[\left(\begin{matrix}8 \\ 8\end{matrix}\right)\left(\begin{matrix}12 \\ 0\end{matrix}\right)\] both of these combinations have a value of 1. it's not crucial to use this method in this problem, but there are situations which are more complex in which it is very useful to use combinations. just saying...

OpenStudy (anonymous):

1

OpenStudy (anonymous):

BTW, let X=# of freshmen chosen. then X ~ Hypergeometric (n=20, k= 8, r = sample size) and \[P(X=x)=\frac{ \left(\begin{matrix}k \\ x\end{matrix}\right) \left(\begin{matrix}n-k \\ r-x\end{matrix}\right)}{ \left(\begin{matrix}n \\ r\end{matrix}\right) }\text{ } \forall x,\,\, 0 \le x \le k \le r\]

OpenStudy (anonymous):

the answer is B.

OpenStudy (anonymous):

compute \[\left(\begin{matrix}20 \\ 8\end{matrix}\right) = \frac{ 20! }{(8!)( 12!) }\]

OpenStudy (anonymous):

show me please...

OpenStudy (anonymous):

just want to make sure you're getting it.

OpenStudy (anonymous):

20/96= 0.20833333333

OpenStudy (anonymous):

sorry... did you follow my example?

OpenStudy (anonymous):